OpenStudy (anonymous):

convert to rectangular (round to tenths): (2, pi/3)

jimthompson5910 (jim_thompson5910):

(r, theta) = (2, pi/3) r = 2 theta = pi/3

jimthompson5910 (jim_thompson5910):

x = r*cos(theta) x = 2*cos(pi/3) x = ???

jimthompson5910 (jim_thompson5910):

y = r*sin(theta) y = 2*sin(pi/3) y = ???

OpenStudy (anonymous):

(1, 1.7) Thanks!

jimthompson5910 (jim_thompson5910):

that's approximate, but you got it pretty much